A man is putting food for fencing around his rectangular backyard that is 53 feet wide. he uses 190 feet of fencing to enclose his backyard. What is the length of the backyard?
Jane is replacing carpet. She bought 143 square ft of carpet. If the length of carpet is 13 feet what is the width?

Answer by Cromlix(4381)

Hi there,
Perimeter of rectangle
= 2 x Width + 2 x Length
190 = 2(53) + 2(x)
190 = 106 + 2x
2x = 190 - 106
2x = 84
x = 42 feet
Length = 42 feet.
..............
Area of carpet = length x width
143 = 13 x width
Width = 143/13
Width = 11 feet
..............
Hope this helps :-)